OpenStudy (anonymous):

what is the value of x 2/7+12/x (use cross products to write the equation:2x=84) Please help!!!!

OpenStudy (firekat97):

do you mean 2/7 = 12/x ?

OpenStudy (anonymous):

yes

OpenStudy (firekat97):

oh okay, so yeah you apply the cross product as you had written up above, to get, 2x = 84 (do you understand this part?)

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

So you do know how to cross multiply ?

OpenStudy (firekat97):

okay now, to find the value of x, we need to isolate it, do you have any ideas on how we can do this?

OpenStudy (texaschic101):

just checking......but do you understand how you got 2x = 84 ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

ok....continue FireKat :)

OpenStudy (firekat97):

okay so we have 2x = 84, so to avoid changing the equation, whatever we do to one side, we must also do to the other, so in this case, to get rid of the coefficient of x, we can divide both sides by 2, so we get; 2x/2 = 84/2 which we can now simplify to x = 42. Do you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the anwser is 42?

OpenStudy (firekat97):

yup because we divided 84 by 2 and that simplifies down to 42

OpenStudy (anonymous):

thank you sososo much

OpenStudy (firekat97):

no problem :)